Circular No. 8120

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003ds AND 2003dt Further to IAUC 8117, J. Graham, W. Li, M. Schwartz, and P. R.

Holvorcem report two apparent supernovae found via LOTOSS on unfiltered KAIT (SN 2003ds) and Tenagra II (SN 2003dt) images:

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003ds Apr. 14.3 10^{h}18^{m}58.^{s}35 +46^{o}27′15′′.5 18.0 4′′.2 E, 1′′.2 N
2003dt Apr. 22.5 20 47 17.56 + 0 18 42.8 17.5 23′′ W, 32′′ S

Additional approximate magnitudes (via KAIT unless otherwise noted): SN 2003ds in NGC 3191, Mar. 25.3 UT, [19.5; Apr. 19.3, 18.5. SN 2003dt in NGC 6962, 2002 Nov. 24.1, [18.8; Apr. 21.5, 17.5 (poor discovery image via Tenagra IV). Schwartz adds that SN 2002ha is at mag ∼ 19.6 on Apr. 22.5, located ∼ 15′′ from SN 2003dt.

SUPERNOVAE 2002km–2002ky Further to IAUC 8119, M. Doi reports the following additional i′ mag-

nitudes (all obtained after subtraction from the reference images; uncertainty 0.1–0.2 mag), all for 2002 Dec. 8: SN 2003km, 23.9; SN 2003kn, 27.3; SN 2003ko, 24.7; SN 2003kp, 25.5; SN 2003kq, 26.4; SN 2003kr, 25.3; SN 2003ks, 26.5; SN 2003kt, 25.1; SN 2003ku, 25.3; SN 2003kv, 26.4; SN 2003kw, 25.4; SN 2003kx, 25.0; SN 2003ky, 25.0.

COMET C/2003 G3 (SOHO) Further to IAUC 8111, D. Hammer reports measurements for a comet

found on C2 and C3 SOHO website images by J. Sachs. Hammer provides the following magnitudes (from C3 images and a clear filter unless noted as C2, which included an orange filter): Apr. 4.113 UT, 7.7; 4.196, 7.6; 4.279, 7.8; 4.363, 7.4; 4.446, 6.9; 4.529, 5.7; 4.564, 4.8 (C2); 4.613, 5.3; 4.621, 4.5 (C2); 4.679, 5.6; 4.685, 4.2 (C2); 4.768, 4.1 (C2); 4.838, 4.6 (C2). The reduced astrometry and parabolic orbital elements (q = 0.017 AU, ω = 175^{o}, Ω = 88^{o}, i = 170^{o}) appear on MPEC 2003-G63.

2003 UT  α_{2000}  δ_{2000}  m_{1}
Apr. 4.029  1^{h}02.^{m}5  + 7^{o}27′  7.5
2003 April 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green